OpenStudy (anonymous):

Solve 3c + 4 = -4c - 17

OpenStudy (sasogeek):

have you tried to solve this?

OpenStudy (anonymous):

Hint: Combine like terms.

OpenStudy (anonymous):

I have , I couldn't do it. Which is why I'm asking you guys. LOL. I'm an English Major. Math is not my cup of tea...

OpenStudy (anonymous):

Combine like terms. Then, solve for c. 3c + 4 = -4c - 17 Subtract 4 from both sides. 3c = -4c - 21 Add 4c to both sides. 7c = -21 Divide both sides by 7. c = -21 /7 So what is c?

OpenStudy (anonymous):

Does that help @GryffindorWins! ?

OpenStudy (anonymous):

Thank you. It does. c:~

OpenStudy (anonymous):

Glad it helped. (: